Evaluate the integrals

i) "\int_{s}^{}\overrightarrow{F} \overrightarrow{ds}"

ii) "\int_{s}^{}\overrightarrow{F} \times \overrightarrow{ds}"


"\overrightarrow{F} = ( x^2 +y , y + z , z + x)" and s is the square 0 ≤ 𝑥 ≤ 1, 0 ≤ 𝑧 ≤ 2 and 𝑦 = 0 positively oriented along the positive Y-axis.


1
Expert's answer
2021-01-28T04:25:06-0500

for this area "\vec{ds}=\vec{j}ds=\vec{j}dxdz"

(i)

"\int_S\vec{F}\vec{ds}=\int_0^2\int_0^1((x^2+y)\vec{i}+(y+z)\vec{j}+(z+x)\vec{k})\vec{j}dxdz="


"=\int_0^2\int_0^1(y+z)dxdz=\int_0^2\int_0^1zdxdz"=

(y=0 on this area)

"=\int_0^2zx|_{x=0}^{x=1}dz=\int_0^2zdz=\frac{z^2}{2}|_0^2=2"


(ii)

"\int_S\vec{F}\times\vec{ds}="


"=\int_S\begin{vmatrix}
\vec{i} & \vec{j} & \vec{k} \\
F_x & F_y & F_z \\
ds_x &ds_y & ds_z
\end{vmatrix}="

(here we multiply with "ds_x=0" and "ds_z=0" )


"=\int_S(\vec{i}(-F_zds_y)+\vec{k}F_xds_y)="


"=\int_0^2\int_0^1(-\vec{i}(z+x)+\vec{k}(x^2+y))dxdz="

(y=0 on this area)

"=\int_0^2\int_0^1(-\vec{i}(z+x)+\vec{k}x^2)dxdz="


"=\int_0^2(-\vec{i}(zx+\frac{x^2}{2})|_{x=0}^{x=1}+\vec{k}\frac{x^3}{3}|_{x=0}^{x=1})dz="


"=\int_0^2(-\vec{i}(z+\frac{1}{2})+\vec{k}\frac{1}{3})dz="


"=-\vec{i}(\frac{z^2}{2}+\frac{z}{2})|_0^2+\vec{k}\frac{z}{3}|_0^2="


"=-3\vec{i}+\frac{2}{3}\vec{k}"